OpenStudy (anonymous):

write an equation of the line containing the given points and perpendicular to the given line. (3,-7); 6x+5y=8

OpenStudy (anonymous):

Why don't you first change that equation in slope-intersect form: \[y=mx+b\]

OpenStudy (anonymous):

*into

OpenStudy (anonymous):

i can get all the way to where you bring y over using the formula y-(-7)=5/6(x-3)

OpenStudy (anonymous):

y+7=5/6x-15/6